# Enhancing Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

### v1

@@ 3.1. Offline Pattern Mining
Output: trees: frequent pattern in the prefix trees form.
1 Function build(ESs):
@@
[…]
             buildTree(S, trees, updated)   // Build trees per each line
3
      return trees   // Embedding trees based on the complete program
4
5Function buildTree(seq, trees, updated):
6       start $\leftarrow$ 0 while start $<$ len(seq) do
             root $\leftarrow$ findOrCreate(trees, seq[start])   // Query existing trees
             root.sup $\leftarrow$ root.sup + 1   // Increase the token frequency
[…]
8Function traverse(tree, seq, skip, length, updated):
9       if length $\geq$ MAX_LEN or len(seq) ==absent$=$ = 0 then
10             return
11       else if skip $<$ MAX_SKIP then
             traverse(tree, seq[1:], skip+1, length)   // Skip the token
12
      node $\leftarrow$ tree.childNode(seq[0])   // Do not skip the token
13       if node is None then
14             node $\leftarrow$ create a new node of id seq[0] tree.setChildNode(seq[0], node)
15       if node is not in updated then
             node.sup $\leftarrow$ node.sup + 1   // Increase token frequency
[…]
17       traverse(node, seq[1:], skip, length + 1)
Algorithm 1 Tree Building and Pattern Construction
According to the preliminary study presented in Section 2.1, most of the reusable code elements for patch generation exist at very fine-grained granularity, i.e., one or two tokens. Therefore, to reuse code elements at the token level, we proposed a new code pattern mining algorithm in Repatt by leveraging the locality property of source code. Specifically, Repatt performs an offline pattern mining process over the given faulty project, which aims to construct a database of token usage patterns that can be used for facilitating the online patch generation process in the subsequent repair process. As aforementioned, we only consider the token patterns within single lines to confine the search space and avoid involving too much noise. In particular, it is common that the tokens used in different code lines share both commonalities and diversities. For example, the token sequence in some code lines can be (“a”, “b”, “c”), while it may also be (“a”, “d”, “c”) in some other code lines. As a result, the sequence (“a”, “c”) is commonly used in both code lines (i.e., commonality) while the tokens “b” and “d” are specific to certain code contexts (i.e., diversity). To preserve the semantics of programs, we keep the token orders and design a skip-fashion pattern mining algorithm to address the problem of diversity among code lines. Also, Repatt embeds each token with a unique id for facilitating the token comparing process. In particular, Repatt removes the separators (e.g., “,” and “(”) and structure keywords (e.g., “if” and “for”) since they are unhelpful for token-level patch generation.
Definition 3.1 ().
(Prefix Embedding Tree): a prefix embedding tree constitutes a set of nodes, each of which is a tuple of $t=\left\langle tok,id,p,C,sup\right\rangle$, where $id$ denotes a unique embedding of a certain token $tok$, $p$ denotes the parent node of $t$, $C$ denotes a set of child nodes of $t$, while $sup$ represents the frequency of the token sequence from the root to the current node $t$.
Furthermore, we use $t.childNode(id)$ to obtain the child node of $t$ with the embedding $id$, and use $t.setChildNode(id,chd)$ to set the node $chd$ as the child node of $t$. According to these notions, we present the pattern mining process in Algorithm 1. In general, given the faulty program, Repatt first decomposes all the code lines in the program into a set of token sequences (i.e., ESs), where each token sequence corresponds to one line of code. Then, it constructs the prefix embedding trees (i.e., trees), which stores the token patterns with corresponding usage frequencies.
Specifically, for each token sequence $S\in$ESs (line 5), Repatt either creates new trees using it or expands existing trees by increasing the frequencies of corresponding patterns (lines 10-11 and 20-26). In particular, we use the function findOrCreate(trees, tok) to find the tree in trees, whose root node corresponds to the token tok, or create a new tree in case it does not exist. Additionally, Repatt employs a skip-fashion pattern mining strategy in this process. That is, for a given token, Repatt will consider both including (lines 20-27) and skipping (lines 18-19) it in the constructed patterns. In particular, we set two hyper-parameters MAX_LEN (line 16) and MAX_SKIP (line 18) to control the max number of tokens in a pattern and the max number of tokens that are allowed to skip. In our evaluation, we will also study their effects on our approach.
Taking the faulty code Codec-3 (Listing 1) as an example, the token sequence of the faulty code line will be $S$=(“contains”, “value”, “index”, “+”, “1”, “4”, “IER”), based on which the constructed prefix embedding trees are presented in Figure 3, where we use the red line to highlight the patterns related to the token sequence $S$. Consequently, the token sequence from the root node to the leaf node denotes a possible pattern and the number denotes its frequency. For example, the frequency of the pattern (“contains”, “value”, “1”, “IER”) is 3, while the pattern (“contains”, “value”, “1”, “3”) is 5. In particular, each distinct token will be the root node of an individual prefix embedding tree. For instance, we present two prefix embedding trees in Figure 3 that are related to the token sequence $S$. As a result, when given a buggy sequence, we can only search a very limited number of embedding trees whose root tokens are included in the sequence, which can promote the searching process. In our pattern mining process, given the minimal frequency of patterns as MIN_SUPPORT, the patterns whose frequency is smaller than MIN_SUPPORT will be deleted. A higher frequency represents a higher possibility of the pattern to be common and reusable for program repair. In the repair process, given the faulty code token sequence $S^{\prime}$, Repatt will search all possible patterns over the embedding trees and generate patches by referencing them.
Figure 3. Partial prefix embedding trees. We use concrete tokens to represent the token ids for ease of understanding.
3.2. Online Code Search and Representation
The expression-level code search aims to find code usage patterns across multi-lines of code for reference, which may cause a large search space if it is also considered like the token-level pattern mining since there may be hundreds or thousands of lines of code even in a single method. As a result, the expression-level pattern mining process is designed as an online code search process, which automatically identifies a set of reference code when given the faulty code. Following existing work (Jiang et al., 2018), Repatt extracts no more than three lines of code respectively before and after the given faulty line as the faulty snippet, and then finds the $N$ most similar snippets via measuring the code similarity between the faulty and reference snippets. Specifically, Repatt extracts a vector from each snippet, where each element in the vector represents a feature and the feature value represents the number of corresponding AST node types in the snippet by following the study (Jiang et al., 2018). Then, we calculate the cosine similarity (Salton, 1988) between two vectors for code ranking.
After obtaining the reference code, existing redundancy-based APR approaches usually try to generate patches by directly referencing the similar code snippets (Yang et al., 2023; Jiang et al., 2018), which is valid if the reference snippets are very similar to the faulty one. However, the contexts of desired reference snippets can be diverse and largely different from the faulty code in real practice, where existing methods will be less effective. To address this issue, we propose a new code representation method, named “Simplified Three-Address Code” (abbr. S-TAC). It unifies the code representations by ignoring the code structures under different contexts but still preserves the locality property of source code.
Definition 3.2 ().
Additionally, we use $T:=t_{1},t_{2}$ to represent $\left\langle T,t_{1},t_{2}\right\rangle$ when there is no ambiguity. Compared with traditional Three-Address Code (TAC) (Aho et al., 2020), S-TAC ignores both the coarse-grained code structure information, e.g., if, and the fine-grained operators, e.g., “+”. The reason for ignoring the structure information is to make the source code from different contexts applicable for patch generation. For example, the conditional expression $a>b$ in if statements can match that either in for statements or in trinary conditional expressions, enabling a more flexible code match. Similarly, the fine-grained operators may also be specific to certain contexts and thus affect the matching and reuse of code. In other words, our S-TAC representation can better reflect the locality property of source code while reducing the noise induced by different contexts.
Taking the code shown in Listing 2 and 3 as examples, the S-TAC form for the if condition from the faulty code is $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.STRING$, while it will be $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.NULL$ for the reference code, where $T_{1}$ and $T_{2}$ are intermediate symbols. In this way, the two conditions can be better matched with each other. We will also evaluate the contribution of our S-TAC form in the experiment (Section 5).
3.3. Reference Code Adaptation
When given the faulty code, we can obtain a set of reference code based on the previous two processes. Then for each reference code pattern, Repatt tries to generate candidate patches by matching the faulty code to the reference code. Algorithm 2 presents the details of the matching process. In general, Repatt performs a greedy match between the faulty code and the reference code. Then, it generates patches according to the difference between them. In other words, given the token (or S-TAC) sequences for both faulty (i.e., BS) and reference code (i.e., RS), the matching algorithm will return a set of pairs (i.e., PS), based on which Repatt generates candidate patches according to a set of pre-defined rules.
[…]
@@ 1 Function matchElement(BS, RS):
       PS [$\leftarrow$[]← [ ], bf $\leftarrow-1$, rf $\leftarrow-1$ cs $\leftarrow LCS(BS,RS)$   // The longest common sequence
2       if  cs.size() $>$ 0 then
@@
4                   bf $\leftarrow BS.indexOf(c.fst)$ rf $\leftarrow RS.indexOf(c.snd)$
@@
5
6       PS normal-$\leftarrow$ PS $\cup$ tryToMatchParent(PS) return PS
7Function tryToMatchParent(PS):
8       result [$\leftarrow$[]← [ ] foreach $\left\langle a,b\right\rangle$ in PS do
9             parent $\leftarrow a.getParent()$ children $\leftarrow parent.getChildren()$ if PS.FirstSet().containsAll(children) then
10                   ts [$\leftarrow$[]← [ ] foreach c in children do
@@ 6       PS normal-$\leftarrow$ PS $\cup$ tryToMatchParent(PS) return PS
Figure 4. Matching results of partial S-TACs in Gson-17 shown in Listing 2. $b_{i}$ represents the S-TAC form of the faulty if statement while the $p_{i}$ corresponds to the if statement in the reference code.
@@
Finally, according to the constructed mapping (i.e., PS), Repatt will generate candidate patches from each pair $(a,b)\in PS$ according to the following rules (we use $a$ to represent the associated AST node in what follows for ease of presentation): (1) replace $a$ with $b$ if the value type of $b$ is compatible with that of $a$; (2) insert $b$ before and after $a$ if $b$ is a standalone statement, e.g, expression statement; (3) insert $b$ before $a$ if $b$ is a conditional expression. In particular, we disable delete operations for patch generation since they tend to generate incorrect patches (Tan et al., 2016; Jiang et al., 2018). Moreover, during this process, Repatt will automatically check the validity of the new code $b$ via static analysis for efficiency, such as whether the variables used by $b$ are all defined in the faulty location. Taking the mapping shown in Figure 4 as an example, one candidate modification will be replacing the expression in.peek()!=JsonToken.STRNG with in.peek()==JsonToken.NULL according to the mapping $(b2,p2)$.
3.4. Patch Ranking
@@
After generating candidate patches, they will be evaluated against the test suites associated with the faulty project. To make the correct patches be evaluated as early as possible and reduce incorrect patches, we have proposed a hierarchical patch ranking strategy. As demonstrated in our preliminary study, most of the reusable code elements are at the fine-grained granularity. Furthermore, previous studies (Mechtaev et al., 2015) also show that correct patches tend to involve small code changes. Therefore, Repatt ranks all the patches generated by the finer-grained token-level code changes higher than those generated by the coarse-grained expression-level code changes. In other words, the patches generated by token-level code changes will be evaluated first and then those by expression-level code changes. Please note that the online pattern mining process introduced in Section 3.1 can be proceeded independently ahead of time and the constructed patterns can be reused for repairing different bugs on demand.
@@
Then, regarding the patches generated by expression-level code changes, Repatt simply takes the similarity in code search (refer to Section 3.2) as the ranking score by following previous studies (Jiang et al., 2018). In contrast, for patches generated by token-level code changes, we consider two important factors: the frequency of the reference code pattern and the similarity between the faulty and fixed code by following existing studies (Mechtaev et al., 2015; Jiang et al., 2018). The ranking score is defined by Formula 1,where $freq$ denotes the frequency of a pattern and $max\_freq$ denotes the max $freq$ of patterns referenced for generating patches. $L_{orig}$ and $L_{fixed}$ represent the length of source code (i.e., number of tokens) before and after applying a patch, while LevenshteinDist represents the Levenshtein distance (Levenshtein, 1966) between the faulty and the fixed code. As a result, a patch with a higher score will rank higher and thus will be validated earlier by running the test cases in the faulty project.
(1) $\small score=0.5*\frac{freq}{max\_freq}+0.5*(1-\frac{{LevenshteinDist}}{max({L_{orig}},{L_{fixed}})})$
4. Experiment Configuration
@@
  • RQ1: How effective is our approach in repairing real-world bugs?
  • RQ2: Can our approach improve the state-of-the-art APRs?
  • RQ3: What is the contribution of each component in Repatt?
@@ 4.1. Subjects and Baselines
In our evaluation, we employed the widely-used Defects4J benchmark (Just et al., 2014) by following existing studies (Xin and Reiss, 2017b; Xuan et al., 2017; Xiong et al., 2017; Yang et al., 2023; Jiang et al., 2023a; Zhu et al., 2021; Jiang et al., 2018; Liu et al., 2019). In particular, we used both Defects4J v1.2 and v2.0 to demonstrate the generality of our approach. Specifically, v1.2 includes 395 bugs from six large-scale real-world projects and v2.0 includes 440 additional bugs from 12 real-world projects.
@@
Furthermore, to show the effectiveness of Repatt, we compared it with eight state-of-the-art APR approaches from different categories, i.e., SimFix (Jiang et al., 2018), TransplantFix (Yang et al., 2023), TBar (Liu et al., 2019), Recoder (Zhu et al., 2021), SelfRepair (Ye et al., 2022b), CodeT5 (Jiang et al., 2023a), CodeGen (Jiang et al., 2023a) and PLBART (Jiang et al., 2023a). Specifically, SimFix and TransplantFix are the two latest and representative redundancy-based methods that repair bugs by referencing similar code. TBar is the best-performing template-based repair technique that generates candidate patches by a set of manually defined patch templates. Finally, Recoder and SelfRepair are the latest and best-performing deep learning methods that are specially designed for program repair, while CodeT5, CodeGen, and PLBART represent the most recent advance in APR techniques by adopting large-scale pre-trained language models. These baselines are the best-performing APR approaches using different technologies and their complete experimental results are available. By comparing our approach with these diverse baselines, we would like to analyze the overall effectiveness of our approach from different perspectives and make the conclusions reliable.
4.2. Configuration and Metrics

### v3

@@ III-A Offline Pattern Mining
Output: trees: frequent pattern in the form of prefix tree.
2
3
4
5Function build(ESs):
6  trees $\leftarrow\emptyset$
7  foreach $S$ in ESs do
     updated $\leftarrow\emptyset$ // Avoid counting duplicated tokens
     buildTree(S, trees, updated) // Build trees
8
 return trees // Embedding trees in programs
9
10Function buildTree(seq, trees, updated):
11  start $\leftarrow$ 0
12  while start << len(seq) do
     root $\leftarrow$ findOrCreate(trees, seq[start]) // Find node
     root.sup $\leftarrow$ root.sup + 1 // Increase the frequency
13     start $\leftarrow$ start + 1
14     traverse(root, seq[start:], 0, 1, updated)
15     trees $\leftarrow$ trees $\cup$ {root}
16
17Function traverse(tree, seq, skip, length, updated):
18  if length $\geq$ MAX_LEN or len(seq) ==== 0 then
19     return
20  else if skip << MAX_SKIP then
21     traverse(tree, seq[1:], skip+1, length, updated)
 node $\leftarrow$ tree.childNode(seq[0]) // Not skip the token
22  if node is None then
23     node $\leftarrow$ create a new node of id seq[0]
24     tree.setChildNode(seq[0], node)
25
26  if node is not in updated then
     node.sup $\leftarrow$ node.sup + 1 // Increase frequency
27     updated $\leftarrow$ updated $\cup$ {node}
28
29  traverse(node, seq[1:], skip, length + 1, updated)
30
Algorithm 1 Tree Building and Pattern Construction
Specifically, for each token sequence $S\in$ ESs (line 3), Repatt either creates new trees using it or expands existing trees by updating pattern frequencies via buildTree(*) (Line 5). In this process, Repatt maintains a set updated to record the nodes whose frequencies have been calculated for avoiding duplicated counting. More specifically, given the token sequence seq and trees that have already constructed, Repatt performs a top-down tree building process, iterating over sub-sequences starting at different positions in seq (Line 9). That is, for each token sub-sequence starting at start, Repatt finds the tree (or creates a new one) whose root node is associated with the token seq[start] (Line 10) and updates its frequency (Line 11). The tree is then recursively expanded via traverse() using a skip-fashion pattern mining strategy, where a token can either be included (Lines 20-27) or skipped (Lines 18-19), constrained by MAX_SKIP (Line 18). For each token, Repatt locates an existing node (Line 20) or creates a new one if absent (Lines 21-23), updating its frequency as needed (Lines 24-26). A pattern is finalized upon reaching the maximum sequence length MAX_LEN. Our evaluation further examines the impact of MAX_LEN and MAX_SKIP on Repatt’s performance.
Figure 3: Partial prefix embedding trees. We use concrete tokens to represent the token ids for ease of understanding.
Taking the faulty code Codec-3 (Listing 1) as an example, the token sequence of the faulty code line will be SS=(“contains”, “value”, “index”, “+”, “1”, “4”, “IER”), based on which the constructed prefix embedding trees are presented in Figure 3, where we use the red line to highlight the patterns related to the token sequence $S$. Consequently, the token sequence from the root node to the leaf node denotes a possible pattern and the number denotes its frequency. For example, the frequency of the pattern (“contains”, “value”, “1”, “IER”) is 3. In particular, each distinct token will be the root node of an individual prefix embedding tree. For instance, we present two prefix embedding trees in Figure 3 that are related to the token sequence $S$. As a result, when given a buggy sequence, we can only search a very limited number of embedding trees whose root tokens are included in the sequence, which can promote the searching process. During pattern mining, patterns with a frequency below a predefined threshold MIN_SUPPORT are discarded. A higher frequency suggests that a pattern is more common and thus more likely to be reusable for repair. In the repair phase, given a faulty code token sequence $S^{\prime}$ , Repatt searches for all matching patterns in the embedding trees and generates patches based on these references.
III-B Online Code Search and Representation
The expression-level code search aims to find code usage patterns across multi-lines of code for reference, which may cause a large search space if it is also considered like the token-level pattern mining since there may be hundreds or thousands of lines of code even in a single method. As a result, the expression-level pattern mining process is designed as an online code search process, which automatically identifies a set of reference code when given the faulty code. Following existing work [8], Repatt extracts no more than three lines of code respectively before and after the given faulty line as the faulty snippet, and then finds the $N$ most similar snippets via measuring the code similarity between the faulty and reference snippets. Specifically, Repatt extracts a vector from each snippet, where each element in the vector represents a feature and the feature value represents the number of corresponding AST node types in the snippet by following the study [8]. Then, we calculate the cosine similarity [49] between two vectors for code ranking.
To adapt the reference code to the buggy code that many have complex contexts in real practice, we propose a new code representation method, named “Simplified Three-Address Code” (abbr. S-TAC). It unifies the code representations by ignoring the code structures under different contexts but still preserves the locality property of source code.
Definition 2
Additionally, we use $T:=t_{1},t_{2}$ to represent $\left\langle T,t_{1},t_{2}\right\rangle$ when there is no ambiguity. Compared with traditional Three-Address Code (TAC) [50], S-TAC ignores both the coarse-grained code structure information, e.g., if, and the fine-grained operators, e.g., “+”. The reason for ignoring the structure information is to make the source code from different contexts applicable for patch generation. For example, the conditional expression $a>b$ in if statements can match that either in for statements or in trinary conditional expressions, enabling a more flexible code match. Similarly, the fine-grained operators may also be specific to certain contexts and thus affect the matching and reuse of code. In other words, our S-TAC representation can better reflect the locality property of source code while reducing the noise induced by different contexts.
Taking the code shown in Listing 2 and 3 as examples, the S-TAC form for the if condition from the faulty code is $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.STRING$, while it will be $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.NULL$ for the reference code, where $T_{1}$ and $T_{2}$ are intermediate symbols. In this way, the two conditions can be better matched with each other. We will also evaluate the contribution of our S-TAC form in the experiment (Section V).
III-C Reference Code Adaptation
When given the faulty code, we can obtain a set of reference code based on the previous two processes. Then for each reference code pattern, Repatt tries to generate candidate patches by matching the faulty code to the reference code. Algorithm 2 presents the details of the matching process. In general, Repatt performs a greedy match between the faulty code and the reference code. Then, it generates patches according to the difference between them. In other words, given the token (or S-TAC) sequences for both faulty (i.e., BS) and reference code (i.e., RS), the matching algorithm will return a set of pairs (i.e., PS), based on which Repatt generates candidate patches according to a set of predefined rules.
@@ III-C Reference Code Adaptation
1
@@ 4 Function matchElement(BS, RS):
5  PS $\leftarrow[]$, bf $\leftarrow-1$, rf $\leftarrow-1$
  cs $\leftarrow LCS(BS,RS)$ // The longest common sequence
6  if  cs.size() >> 0 then
@@ 5  PS $\leftarrow[]$, bf $\leftarrow-1$, rf $\leftarrow-1$
8        os $\leftarrow BS.getNodes(bf,BS.indexOf(c.fst))$
9        ts $\leftarrow RS.getNodes(rf,RS.indexOf(c.snd))$
@@
10        bf $\leftarrow BS.indexOf(c.fst)$
11        rf $\leftarrow RS.indexOf(c.snd)$
12
13    os $\leftarrow BS.getNodes(bf,BS.length)$
14     ts $\leftarrow RS.getNodes(rf,RS.length)$
@@
15
16  PS $\leftarrow$ PS $\cup$ tryToMatchParent(PS)
17  return PS
18Function tryToMatchParent(PS):
19  result $\leftarrow[]$
20  foreach $\left\langle a,b\right\rangle$ in PS do
21     parent $\leftarrow a.getParent()$
22     children $\leftarrow parent.getChildren()$
23     if PS.FirstSet().containsAll(children) then
        // All sibling AST nodes are not matched
24        ts $\leftarrow[]$
25        foreach c in children do
@@ 16  PS $\leftarrow$ PS $\cup$ tryToMatchParent(PS)
Figure 4: Matching results of partial S-TACs in Gson-17 shown in Listing 2. $b_{i}$ represents the S-TAC form of the faulty if statement while the $p_{i}$ corresponds to the if statement in the reference code. In particular, we also present the source code corresponding to each T-SAC.
@@
Finally, according to the constructed mapping (i.e., PS), Repatt will generate candidate patches from each pair $(a,b)\in PS$ according to the following rules (we use $a$ to represent the associated AST node in what follows for ease of presentation): (1) replace $a$ with $b$ if the value type of $b$ is compatible with that of $a$; (2) insert $b$ before and after $a$ if $b$ is a standalone statement, e.g, expression statement; (3) insert $b$ before $a$ if $b$ is a conditional expression. Both token-level and expression-level patch generations follow this mapping algorithm, based on the types of $a$ and $b$ (e.g., replacing variable $a$ with another variable $b$ generates a token-level patch). In particular, we disable delete operations for patch generation since they tend to generate incorrect patches [5, 8]. Moreover, during this process, Repatt incorporates a built-in validation step that performs lightweight static analysis to check the validity of the newly generated code $b$, such as whether all variables used by $b$ are valid and usable in the faulty location by checking their scopes.
III-D Patch Ranking
@@
After generating candidate patches, they will be evaluated against the test suites associated with the faulty project. To make the correct patches be evaluated as early as possible and reduce incorrect patches, we have proposed a hierarchical patch ranking strategy. As demonstrated in our preliminary study, most of the reusable code elements exist at the fine-grained granularity. Furthermore, previous study [51] also shows that correct patches tend to involve small code changes. Therefore, Repatt ranks all the patches generated by the finer-grained token-level code changes higher than those generated by the coarse-grained expression-level code changes.
@@
Then, regarding the patches generated by expression-level code changes, Repatt simply takes the similarity in code search (refer to Section III-B) as the ranking score by following previous studies [8]. For patches generated by token-level code changes, we consider two important factors: the frequency of the reference code pattern and the similarity between the faulty and fixed code by following existing studies [51, 8]. The ranking score is defined by Formula 1,where freqfreq denotes the frequency of a pattern and max_freqmax\_freq denotes the max freqfreq of patterns referenced for generating patches. $L_{orig}$ and $L_{fixed}$ represent the length of source code (i.e., number of tokens) before and after applying a patch, while LevenshteinDist represents the token-level Levenshtein distance [52] between the faulty and the fixed code. As a result, a patch with a higher score (i.e., referencing a more frequent pattern and applying finer-grained modifications) will rank higher and thus will be validated earlier by running the test cases in the faulty project.
$\small score=0.5*\frac{freq}{max\_freq}+0.5*(1-\frac{{LevenshteinDist}}{max({L_{orig}},{L_{fixed}})})$ (1)
IV Experiment Configuration
@@
  • RQ1: How effective is Repatt in repairing real-world bugs , compared to the state-of-the-art APRs?
  • RQ2: Can a combination of Repatt with traditional APR approaches improve the state-of-the-art APRs?
  • RQ3: What is each component’s contribution in Repatt?
@@ IV-A Subjects and Baselines
In our evaluation, we employed the widely-used Defects4J benchmark [41], following existing studies [9, 3, 2, 33, 20, 18, 8, 4]. In particular, we used both Defects4J v1.2 and v2.0 to demonstrate the generality of our approach. Specifically, v1.2 includes 395 bugs from six large-scale real-world projects and v2.0 includes 440 additional bugs from 12 real-world projects.
@@
Furthermore, to show the effectiveness of Repatt, we compared it with ten state-of-the-art APR approaches from different categories, i.e., CapGen [1], SimFix [8], TransplantFix [33], TBar [4], Recoder [18], SelfAPR [19], ITER [53], AlphaRepair [22], Repilot [23] and GAMMA [24]. Specifically, CapGen, SimFix and TransplantFix are the three latest and representative redundancy-based methods that repair bugs by referencing similar code. TBar is the best-performing template-based repair technique that generates candidate patches by a set of manually defined patch templates. Recoder, ITER and SelfAPR are the latest and best-performing deep learning methods that are specially designed for program repair. Finally, AlphaRepair, Repilot, and GAMMA represent the most recent advance in APR techniques by adopting LLMs. These baselines are the best-performing APR approaches using different technologies and their complete experimental results are available. By comparing our approach with these diverse baselines, we would like to analyze the overall effectiveness of our approach from different perspectives and make the conclusions reliable.
IV-B Configuration and Metrics